Swarm Operational News
19 September 2017
The Department of Physics and Astronomy at the University of Iowa seeks to appoint a post-doctoral researcher to join a dynamic and expanding group of space physics researchers. Amongst other sources, the post-holder will be using Swarm mission data to aid in studies of the Earth's magnetic field.
18 September 2017
We are pleased to announce a new release of non-gravitational acceleration and thermosphere neutral density data, which was derived from the measurements of the GPS receiver.
14 September 2017
We inform the Swarm users that due to an anomaly that occurred on the VFM (Vector Field Magnetometer) on board Swarm Bravo, i.e., VFM generation time of science packet frozen from 11/09/2017 03:54 UTC to 13/09/2017 12:24 UTC, the Swarm Level 0 data for this time window was not generated.
13 September 2017
With reference to the news published on 08 July 2016 we inform the Swarm users that the ASM-VFM Residual dataset produced for the investigations of the scalar residuals between the readings of the scalar magnetometer (ASM) and the vector magnetometer (VFM) has been extended until 31 August 2017 for all Swarm spacecraft.
13 September 2017
We are pleased to inform the Swarm users that the 4-days gap present in the Swarm Alpha MAGx_LR_1B data almost at the beginning of the mission, i.e., from the 14 to 17 December 2013, has been recovered.
07 September 2017
We are pleased to inform the Swarm users that an improved version of the MAGx_LR_1B products (Product Baseline and File Counter 0502) is now available in the "Advanced" folder of the ESA FTP server. This dataset covers the time window from the beginning of the mission to end August 2017 for all Swarm spacecraft.
We would like to inform the Swarm users that an Orbit Maintenance Manoeuvre was successfully performed on Swarm Alpha on 13 July 2017 starting at 14:17:55 UTC.
In relation to the previous communication we are pleased to inform you that the announced changes to the FTP layout are now effective as planned.
We inform the Swarm users that, starting from 05 July 2017, the Absolute Vertical TEC (VTEC) values and corresponding elevation angles will be available in the Swarm Level 2 Cat-2 TEC products.
We inform the Swarm Users that because of an issue during the retrieval of the telemetry data from Swarm Charlie, a data gap is present in the Level 1B production on 04 July 2017 from 06:02 to 13:43 UTC.
Following the previous communication, we are pleased to inform you that the announced changes to the FTP layout will be effective on 12 July 2017. Please be aware that some service disruptions might happen during that day.
27 June 2017
We are pleased to inform that, following a Swarm community recommendation, several changes will be introduced in the Swarm Dissemination (swarm-diss.eo.esa.int) FTP structure in early July 2017.
19 June 2017
We inform the Swarm users that because of an issue during the processing of the Swarm L1B Orbit and Attitude product of 03/06/2017 on Swarm Alpha and on 08/06/2017 on Swarm Bravo - i.e., available Level 0 data was not included in in the L1B processing - the L1B data production has been affected.
Due to an anomaly that occurred on the VFM (Vector Field Magnetometer) on board Swarm Alpha, the VFM science data (at 50 Hz rate) was not received from 29/04/2017 11:22 UTC to 01/05/2017 08:05 UTC.
27 April 2017
The third release of the Swarm magnetic field models is now available to users. The new models are based on improved Swarm magnetometer data (MAGx_LR_1B data products, baseline 05) covering the period from launch to the end of 2016, i.e. one additional year in comparison to the previous release.
This announcement is to solicit further ideas for new data products and services supporting the outcome of the European Space Agency's Swarm constellation mission.
19 April 2017
Swarm DISC has now, upon request, granted extension of deadline for the following Invitations to Tender:
Swarm DISC ITT 1.1: Swarm as a gravity field mission
Swarm DISC ITT 1.2: Strength and location of the auroral electrojets based on Swarm data
Due to a technical problem in the production chain, Swarm Charlie Level 1b data of 16 March 2017 are currently not available. Consequently, the Swarm Charlie Level 2 Cat-2 (FAC, IBI and TEC) are also not available for this day.
04 April 2017
Swarm DISC has today granted an extension of deadline to the Invitations to Tender SD-ITT-1.1: Swarm DISC ITT 1.1: "Swarm as a gravity field mission" (100,000 EUR, New Closing Date: 09-05-2017).
We inform the Swarm users that an experimental Swarm EFI 2 Hz TII cross-track ion flow dataset is now available in the /Advanced/Plasma_Data/2Hz_TII_Cross-track_Dataset folder on the Swarm ftp server. Data cover from 01/01/2016 to 31/01/2017 for S/C A and B, Product Baseline and File Counter 0101.
Showing 1 - 20 of 132 results.